OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Choose the correct ordered pair. y = 8x - 4 x = 4

OpenStudy (anonymous):

y= 8x-4 y= 8(4)-4 y= 32-4 y= 18 Is the ordered pair (4,18)? owo

OpenStudy (anonymous):

no its not...but thanks!